# Ar2Can: An Architect and an Artist Leveraging a Canvas for Multi-Human Generation

### v1

@@ 1 Introduction
Text-to-image diffusion models [rombach2022high, podell2023sdxl, flux2024] have achieved photorealistic synthesis quality. However, they systematically fail on a fundamental task: generating scenes with multiple distinct humans [borse2025multihuman]. Given a prompt such as “five friends having coffee” and reference identity images, current approaches produce images with duplicated faces, merged identities, or incorrect person counts (Fig. 6). This failure persists across different paradigms: regional conditioning methods [li2023gligen, xu2025withanyone, zhang2025id] require explicit spatial annotations at inference time, limiting usability; identity-preserving approaches [guo2024pulid, labs2025flux1kontext] excel at single-subject personalization but struggle with multi-identity composition; and even methods explicitly designed for multi-ID generation [wu2025omnigen2, xiao2025omnigen, mou2025dreamo, chen2025xverse, cheng2025umo] fail on recently introduced benchmarks [borse2025multihuman] (Section 4). This widespread failure reflects a fundamental limitation in how models represent and compose multiple identities within a shared scene.
@@
Our Approach. We propose Ar2Can (Architect, Artist and Canvas), a two-stage framework that disentangles spatial planning from identity rendering. As illustrated in Fig. 1, the Architect first generates a structured spatial plan–bounding boxes (and optionally pose) specifying where(and how) each person should appear. We design two Architect variants with complementary trade-offs: Architect-A uses Qwen [qwen2024] with supervised fine-tuning on bounding box prediction (L1 + GIoU losses), providing structured spatial reasoning; Architect-B uses Flux-Schnell [flux2024] finetuned with reinforcement learning, enabling adaptive planning with human pose.
The Artist then renders the photorealistic image conditioned on this spatial plan. We train a single Artist model, Flux-Kontext, via Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] that works with either Architect variant. Critically, we introduce a spatially-grounded face matching reward that operates in two stages: (1) Hungarian algorithm matching between expected centroids (from the Architect) and detected face locations, and (2) ArcFace [deng2019arcface] identity similarity between reference faces and their Hungarian-matched generated counterparts. Combined with count accuracy, pose correction and HPSv3 quality rewards [ma2025hpsv3], this compositional reward structure ensures faces appear at correct locations and match the right identities.
@@
    Our evaluations on MultiHuman-Testbench [borse2025multihuman] show substantial improvements in count accuracy and identity preservation over state-of-the-art methods while maintaining competitive perceptual quality and action scores.
@@ 2 Related Work
Text-to-Image Diffusion Models. Recent advances in diffusion models [rombach2022high, saharia2022photorealistic, ramesh2022hierarchical, podell2023sdxl] have achieved remarkable realism. However, these models struggle with compositional tasks [chefer2023attend, liu2022compositional], particularly multi-object and multi-human generation. Flow-based models [esser2024scaling, flux2024] offer faster sampling but inherit similar compositional limitations [disco2025].
Multi-Object and Regional Control. GLIGEN [li2023gligen], ReCo [yang2023reco], and LayoutDiffusion [zheng2023layoutdiffusion] inject spatial control via bounding boxes or layouts at inference time. BoxDiff [xie2023boxdiff] similarly conditions on box annotations. While effective, these methods require explicit spatial annotations from users, limiting practical deployment. Our Architect automates layout generation from text prompts alone.
Identity-Preserving Generation. IP-Adapter [ye2023ip] and InstantID [wang2024instantid] enable training-free identity injection. Methods like PuLID [guo2024pulid] enhance identity fidelity with multi-view consistency. However, these approaches focus on single-subject personalization and struggle when composing multiple identities.
Multi-Human Generation. Recent approaches span identity-preserving diffusion [labs2025flux1kontext], regional conditioning [xu2025withanyone, zhang2025id], and unified multimodal architectures [wu2025omnigen2, xiao2025omnigen, mou2025dreamo, chen2025xverse, cheng2025umo]. MultiHuman-Testbench [borse2025multihuman] exposes systematic failures across all methods: widespread identity duplication, motivating our spatially-grounded approach.
Reinforcement Learning for Diffusion. DDPO [black2023training] introduced RL fine-tuning for diffusion models using policy gradients. ImageReward [xu2023imagereward] and HPS [ma2025hpsv3] provide learned reward functions for human preference alignment. DPOK [fan2023dpok] and AlignProp [prabhudesai2023aligning] extend RL to various objectives. GRPO [shao2024deepseekmath], introduced for language model reasoning, computes group-normalized advantages, reducing variance. DisCo [disco2025] recently applied GRPO to multi-human generation with diversity-focused rewards. Our work extends this paradigm with spatially-grounded rewards that explicitly align face locations with planned layouts.
@@ 3.1 Overview
Ar2Can decomposes multi-human image generation into two stages (Fig. 1). Given a text prompt $p$ (e.g., “three people playing basketball”) and $N$ reference identity images $\{I_{\text{ref},1},\ldots,I_{\text{ref},N}\}$, the Architect module $\psi$ predicts a spatial layout $\mathcal{L}=\{b_{1},\ldots,b_{N}\}$, where each bounding box $b_{i}=(x_{i},y_{i},w_{i},h_{i})$ defines a region with centroid $c_{i}=(x_{i}+w_{i}/2,y_{i}+h_{i}/2)$. Conditioned on the prompt, layout, and reference images, the Artist module $\pi_{\theta}$ synthesizes the final image via $x\sim\pi_{\theta}(x\mid p,\mathcal{L},\{I_{\text{ref}}\})$. This decomposition offers two key advantages: (1) Explicit spatial grounding: The face locations are determined prior to rendering, reducing spatial conflicts that can lead to identity merging. (2) Modular design: the Architect module can be replaced based on deployment constraints (e.g., speed vs. accuracy) without requiring retraining of the Artist. This enables us to benchmark methods such as WithAnyone [xu2025withanyone] and ID-Patch [zhang2025id] using our proposed architects.
@@ 3.2 Training Data Curation
A central challenge in multi-human image generation is the limited availability of supervised training data. Ideally, one would fine-tune models on millions of images featuring multiple individuals within the same scene, each accompanied by distinct reference images representing their identities. Such large-scale datasets are typically proprietary, confined to organizations with access to vast user-generated content(such as Snap or Meta). Public datasets for Multi-Human generation(such as LAION-Face, Multi-ID-2M [xu2025withanyone]) contain lots of images sampled from videos, which aren’t high quality. There are very few images of “Portrait-style, High-quality” open source data; hence methods such as [zhang2025id] curate datasets by purchasing data. However, rather than relying on this scarce real multi-person data, we leverage DisCo [disco2025] to generate synthetic multi-person scenes, then construct hybrid training samples by pairing these scenes with real reference faces. Instead of supervising our Artist with unavailable data, we reinforce it’s predictions using this curated data. This grounds the model in authentic facial appearances and capitalizes on DisCo [disco2025]’s compositional capabilities.
@@
    Multi-view references ($\sim$100K identities): Public datasets consisting of individuals with an additional view from a different angle [guo2016ms, xu2025withanyone], enabling multi-view identity learning.
@@
    Single-view references ($\sim$500K images): Public datasets with single high-quality images lacking a second view [yang2024consistentid]. For a percentage of these, we synthetically generate secondary views using PuLID [guo2024pulid] to introduce pose and expression diversity. For the rest, we apply random augmentations including rotations, horizontal flipping, and brightness adjustment.
@@
    Synthetic faces: Identity-varied faces generated during DisCo [disco2025] scene synthesis. These enhance robustness through synthetic diversity.
@@ 3.3.1 Architect-A: Bounding-Box Coordinate Regression
Directly regressing bounding-box coordinates from text prompts is suboptimal, as it faces two challenges: (i) the number of boxes varies dynamically with the prompt, and (ii) extracting counts requires deep linguistic reasoning (e.g., “Two girls next to a man and behind a lady”). We address this issue by building Architect-A as an autoregressive model using Qwen-2.5 (0.5B) [qwen2024], which provides both dynamic output length and inherent language understanding.
Figure 3 illustrates our Architect-A training pipeline. As observed, we extend the tokenizer with layout structure tokens: <SoL> (start of layout), <EoL> (end of layout) and <C> (coordinate placeholder). Since coordinates require continuous values that standard LLMs cannot directly output, we augment the model with additional prediction heads. Following prior work [lisa, chatpose, aipparel, chatgarment], we attach two heads to the base LLM: a coordinate regression head $f_{\text{value}}$ and a coordinate embedding head $f_{\text{embed}}$, alongside the original token head $f_{\text{token}}$.
@@
where $b_{\text{pred}}=f_{\text{value}}(h_{\text{prev}})$ are predicted coordinates from LLM hidden states $h_{\text{prev}}$, $y_{\text{next}}$ denotes the next token to be predicted, and $b_{\text{gt}}$ are ground-truth boxes. We use gIoU [rezatofighi2019generalized] to handle non-overlapping boxes and L1 loss for stable gradients with normalized coordinates. We also observed that MLP layers are highly sensitive to permutation of bounding box coordinate embeddings, negatively affecting performance and convergence. Hence, we enforce consistent coordinate ordering via data sorting. This simple strategy improves training stability and enables more reliable coordinate regression. Additional implementation details and explanations are presented in the supplementary material.
@@ 3.3.2 Architect-B: T2I-based Layout Generation
We adopt Flux-Schnell [flux2024] as the base T2I model for layout sketching. It is a widely used flow-matching model that achieves high speed by requiring only 4 denoising steps. However, it struggles with accurate person count generation and action alignment in multi-human scenes (see Appendix D). To address these limitations, we fine-tune Flux-Schnell using Group Relative Policy Optimization (GRPO), using a setup similar to DisCo [disco2025]. It recently demonstrated an effective RL Framework for improving count accuracy (and diversity) in multi-human generation.
@@
where $r_{\text{count}}(x)=\mathbbm{1}[n_{\text{pred}}(x)=n_{\text{target}}]$ rewards exact person count matching (detected via blob analysis), and $r_{\text{hps}}(x,p)=\text{HPSv3}(x,p)$ measures prompt alignment using Human Preference Score v3 [ma2025hpsv3]. The fine-tuned Schnell model generates both face bounding boxes and human pose coordinates for downstream use by the Artist. Training hyperparameters are detailed in Appendix B.
@@ 3.4.1 GRPO Training with Compositional Rewards
To address this pressing issue, we propose a two-stage matching strategy. This is illustrated in Figure 4. Given the Architect’s predicted bounding boxes $\{b_{1}^{\text{pred}},\ldots,b_{N}^{\text{pred}}\}$ with centroids $\{c_{i}^{\text{pred}}\}$ and RetinaFace [deng2020retinaface] detected boxes $\{b_{1}^{\text{det}},\ldots,b_{O}^{\text{det}}\}$ with centroids $\{c_{j}^{\text{det}}\}$, we first establish spatial correspondence via Hungarian matching [kuhn1955hungarian] on centroid distance cost matrix $C_{ij}=\|c_{i}^{\text{pred}}-c_{j}^{\text{det}}\|_{2}$, finding optimal assignment $\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{\min(N,O)}C_{i,\pi(i)}$. This centroid-based matching eases the constraint on exact localization, i.e. the Artist must generate faces near the specified locations rather than at precise coordinates, enabling photorealistic rendering with natural variations in pose, scale, and positioning.
For each reference image $I_{\text{ref},i}$, we then compute identity similarity with its spatially-matched generated face using ArcFace [deng2019arcface] embeddings $e_{i}^{\text{ref}}=f_{\text{ArcFace}}(I_{\text{ref},i})$ and $e_{\pi^{*}(i)}^{\text{gen}}=f_{\text{ArcFace}}(\text{crop}(x,b_{\pi^{*}(i)}^{\text{det}}))$:
@@
Pose Rewards. To further reduce copy-paste artifacts and push for natural looking scenes, we compute a frontal pose reward, with the prompt “Everyone is looking at the camera”. Following geometric head pose estimation methods [wu2017simultaneous], we use 5-point facial landmarks to compute a score $\delta_{i}$ based on roll angle and facial symmetry for each detected face $i$. The pose reward is formulated as $r_{\text{pose}}(x,\mathcal{L})=\frac{1}{O}\sum_{i=1}^{O}\delta_{i}$. We provide the formulation for $\delta_{i}$ in Appendix C. Additionally, we also explore fine-grained body pose control paired with with pose canvases, detailed in Appendix C.
@@ 4.1 Experimental Setup
Datasets. We evaluate on two benchmarks: MultiHuman-Testbench (MHTB) [borse2025multihuman] with 1,800 samples and 5550 faces, uniformly spanning 1-5 people per sample; and Multi-ID-Test [xu2025withanyone] with 1-4 faces biased toward 1-2 person scenes.
Evaluation Metrics. Following the setup in MHTB [borse2025multihuman], we employ four complementary metrics: Count Accuracy for correct number of generated people; Hungarian ID Similarity evaluates identity preservation using facial similarity with Hungarian matching to establish and quantify optimal face correspondences; HPSv2 measures overall image quality and prompt alignment; and Action-S/C probes action correctness separately for simple and complex actions, using Gemini Flash as the MLLM. We also report the Unified Metric [borse2025multihuman] which balances identity preservation and prompt alignment.

### v2

@@ 1 Introduction
Text-to-image diffusion models [29, 25, 18] have achieved photorealistic synthesis quality. However, they systematically fail on a fundamental task: generating scenes with multiple distinct humans [4]. Given a prompt such as “five friends having coffee” and reference identity images, current approaches produce images with duplicated faces, merged identities, or incorrect person counts (Fig. 6). This failure persists across different paradigms: regional conditioning methods [20, 38, 42] require explicit spatial annotations at inference time, limiting usability; identity-preserving approaches [15, 2] excel at single-subject personalization but struggle with multi-identity composition; and even methods explicitly designed for multi-ID generation [34, 36, 23, 7, 8] fail on recently introduced benchmarks [4] (Section 4). This widespread failure reflects a fundamental limitation in how models represent and compose multiple identities within a shared scene.
@@
Our Approach. We propose Ar2Can (Architect, Artist and Canvas), a two-stage framework that disentangles spatial planning from identity rendering. As illustrated in Fig. 1, the Architect first generates a structured spatial plan–bounding boxes (and optionally pose) specifying where(and how) each person should appear. We design two Architect variants with complementary trade-offs: Architect-A uses Qwen [32] with supervised fine-tuning on bounding box prediction (L1 + GIoU losses), providing structured spatial reasoning; Architect-B uses Flux-Schnell [18] finetuned with reinforcement learning, enabling adaptive planning with human pose.
The Artist then renders the photorealistic image conditioned on this spatial plan. We train a single Artist model, Flux-Kontext, via Group Relative Policy Optimization (GRPO) [31] that works with either Architect variant. Critically, we introduce a spatially-grounded face matching reward that operates in two stages: (1) Hungarian algorithm matching between expected centroids (from the Architect) and detected face locations, and (2) ArcFace [10] identity similarity between reference faces and their Hungarian-matched generated counterparts. Combined with count accuracy, pose correction and HPSv3 quality rewards [22], this compositional reward structure ensures faces appear at correct locations and match the right identities.
@@
    Our evaluations on MultiHuman-Testbench [4] show substantial improvements in count accuracy and identity preservation over state-of-the-art methods while maintaining competitive perceptual quality and action scores.
@@ 2 Related Work
Text-to-Image Diffusion Models. Recent advances in diffusion models [29, 30, 27, 25] have achieved remarkable realism. However, these models struggle with compositional tasks [6, 21], particularly multi-object and multi-human generation. Flow-based models [11, 18] offer faster sampling but inherit similar compositional limitations [5].
Multi-Object and Regional Control. GLIGEN [20], ReCo [40], and LayoutDiffusion [43] inject spatial control via bounding boxes or layouts at inference time. BoxDiff [37] similarly conditions on box annotations. While effective, these methods require explicit spatial annotations from users, limiting practical deployment. Our Architect automates layout generation from text prompts alone.
Identity-Preserving Generation. IP-Adapter [41] and InstantID [33] enable training-free identity injection. Methods like PuLID [15] enhance identity fidelity with multi-view consistency. However, these approaches focus on single-subject personalization and struggle when composing multiple identities.
Multi-Human Generation. Recent approaches span identity-preserving diffusion [2], regional conditioning [38, 42], and unified multimodal architectures [34, 36, 23, 7, 8]. MultiHuman-Testbench [4] exposes systematic failures across all methods: widespread identity duplication, motivating our spatially-grounded approach.
Reinforcement Learning for Diffusion. DDPO [3] introduced RL fine-tuning for diffusion models using policy gradients. ImageReward [39] and HPS [22] provide learned reward functions for human preference alignment. DPOK [12] and AlignProp [26] extend RL to various objectives. GRPO [31], introduced for language model reasoning, computes group-normalized advantages, reducing variance. DisCo [5] recently applied GRPO to multi-human generation with diversity-focused rewards. Our work extends this paradigm with spatially-grounded rewards that explicitly align face locations with planned layouts.
@@ 3.1 Overview
Ar2Can decomposes multi-human image generation into two stages (Fig. 1). Given a text prompt $p$ (e.g., “three people playing basketball”) and $N$ reference identity images $\{I_{\text{ref},1},\ldots,I_{\text{ref},N}\}$, the Architect module $\psi$ predicts a spatial layout $\mathcal{L}=\{b_{1},\ldots,b_{N}\}$, where each bounding box $b_{i}=(x_{i},y_{i},w_{i},h_{i})$ defines a region with centroid $c_{i}=(x_{i}+w_{i}/2,y_{i}+h_{i}/2)$. Conditioned on the prompt, layout, and reference images, the Artist module $\pi_{\theta}$ synthesizes the final image via $x\sim\pi_{\theta}(x\mid p,\mathcal{L},\{I_{\text{ref}}\})$. This decomposition offers two key advantages: (1) Explicit spatial grounding: The face locations are determined prior to rendering, reducing spatial conflicts that can lead to identity merging. (2) Modular design: the Architect module can be replaced based on deployment constraints (e.g., speed vs. accuracy) without requiring retraining of the Artist. This enables us to benchmark methods such as WithAnyone [38] and ID-Patch [42] using our proposed architects.
@@ 3.2 Training Data Curation
A central challenge in multi-human image generation is the limited availability of supervised training data. Ideally, one would fine-tune models on millions of images featuring multiple individuals within the same scene, each accompanied by distinct reference images representing their identities. Such large-scale datasets are typically proprietary, confined to organizations with access to vast user-generated content(such as Snap or Meta). Public datasets for Multi-Human generation(such as LAION-Face, Multi-ID-2M [38]) contain lots of images sampled from videos, which aren’t high quality. There are very few images of “Portrait-style, High-quality” open source data; hence methods such as [42] curate datasets by purchasing data. However, rather than relying on this scarce real multi-person data, we leverage DisCo [5] to generate synthetic multi-person scenes, then construct hybrid training samples by pairing these scenes with real reference faces. Instead of supervising our Artist with unavailable data, we reinforce it’s predictions using this curated data. This grounds the model in authentic facial appearances and capitalizes on DisCo [5]’s compositional capabilities.
@@
    Multi-view references ($\sim$100K identities): Public datasets consisting of individuals with an additional view from a different angle [14, 38], enabling multi-view identity learning.
@@
    Single-view references ($\sim$500K images): Public datasets with single high-quality images lacking a second view [16]. For a percentage of these, we synthetically generate secondary views using PuLID [15] to introduce pose and expression diversity. For the rest, we apply random augmentations including rotations, horizontal flipping, and brightness adjustment.
@@
    Synthetic faces: Identity-varied faces generated during DisCo [5] scene synthesis. These enhance robustness through synthetic diversity.
@@ 3.3.1 Architect-A: Bounding-Box Coordinate Regression
Directly regressing bounding-box coordinates from text prompts is suboptimal, as it faces two challenges: (i) the number of boxes varies dynamically with the prompt, and (ii) extracting counts requires deep linguistic reasoning (e.g., “Two girls next to a man and behind a lady”). We address this issue by building Architect-A as an autoregressive model using Qwen-2.5 (0.5B) [32], which provides both dynamic output length and inherent language understanding.
Figure 3 illustrates our Architect-A training pipeline. As observed, we extend the tokenizer with layout structure tokens: <SoL> (start of layout), <EoL> (end of layout) and <C> (coordinate placeholder). Since coordinates require continuous values that standard LLMs cannot directly output, we augment the model with additional prediction heads. Following prior work [19, 13, 24, 1], we attach two heads to the base LLM: a coordinate regression head $f_{\text{value}}$ and a coordinate embedding head $f_{\text{embed}}$, alongside the original token head $f_{\text{token}}$.
@@
where $b_{\text{pred}}=f_{\text{value}}(h_{\text{prev}})$ are predicted coordinates from LLM hidden states $h_{\text{prev}}$, $y_{\text{next}}$ denotes the next token to be predicted, and $b_{\text{gt}}$ are ground-truth boxes. We use gIoU [28] to handle non-overlapping boxes and L1 loss for stable gradients with normalized coordinates. We also observed that MLP layers are highly sensitive to permutation of bounding box coordinate embeddings, negatively affecting performance and convergence. Hence, we enforce consistent coordinate ordering via data sorting. This simple strategy improves training stability and enables more reliable coordinate regression. Additional implementation details and explanations are presented in the supplementary material.
@@ 3.3.2 Architect-B: T2I-based Layout Generation
We adopt Flux-Schnell [18] as the base T2I model for layout sketching. It is a widely used flow-matching model that achieves high speed by requiring only 4 denoising steps. However, it struggles with accurate person count generation and action alignment in multi-human scenes (see Appendix D). To address these limitations, we fine-tune Flux-Schnell using Group Relative Policy Optimization (GRPO), using a setup similar to DisCo [5]. It recently demonstrated an effective RL Framework for improving count accuracy (and diversity) in multi-human generation.
@@
where $r_{\text{count}}(x)=\mathbbm{1}[n_{\text{pred}}(x)=n_{\text{target}}]$ rewards exact person count matching (detected via blob analysis), and $r_{\text{hps}}(x,p)=\text{HPSv3}(x,p)$ measures prompt alignment using Human Preference Score v3 [22]. The fine-tuned Schnell model generates both face bounding boxes and human pose coordinates for downstream use by the Artist. Training hyperparameters are detailed in Appendix B.
@@ 3.4.1 GRPO Training with Compositional Rewards
To address this pressing issue, we propose a two-stage matching strategy. This is illustrated in Figure 4. Given the Architect’s predicted bounding boxes $\{b_{1}^{\text{pred}},\ldots,b_{N}^{\text{pred}}\}$ with centroids $\{c_{i}^{\text{pred}}\}$ and RetinaFace [9] detected boxes $\{b_{1}^{\text{det}},\ldots,b_{O}^{\text{det}}\}$ with centroids $\{c_{j}^{\text{det}}\}$, we first establish spatial correspondence via Hungarian matching [17] on centroid distance cost matrix $C_{ij}=\|c_{i}^{\text{pred}}-c_{j}^{\text{det}}\|_{2}$, finding optimal assignment $\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{\min(N,O)}C_{i,\pi(i)}$. This centroid-based matching eases the constraint on exact localization, i.e. the Artist must generate faces near the specified locations rather than at precise coordinates, enabling photorealistic rendering with natural variations in pose, scale, and positioning.
For each reference image $I_{\text{ref},i}$, we then compute identity similarity with its spatially-matched generated face using ArcFace [10] embeddings $e_{i}^{\text{ref}}=f_{\text{ArcFace}}(I_{\text{ref},i})$ and $e_{\pi^{*}(i)}^{\text{gen}}=f_{\text{ArcFace}}(\text{crop}(x,b_{\pi^{*}(i)}^{\text{det}}))$:
@@
Pose Rewards. To further reduce copy-paste artifacts and push for natural looking scenes, we compute a frontal pose reward, with the prompt “Everyone is looking at the camera”. Following geometric head pose estimation methods [35], we use 5-point facial landmarks to compute a score $\delta_{i}$ based on roll angle and facial symmetry for each detected face $i$. The pose reward is formulated as $r_{\text{pose}}(x,\mathcal{L})=\frac{1}{O}\sum_{i=1}^{O}\delta_{i}$. We provide the formulation for $\delta_{i}$ in Appendix C. Additionally, we also explore fine-grained body pose control paired with with pose canvases, detailed in Appendix C.
@@ 4.1 Experimental Setup
Datasets. We evaluate on two benchmarks: MultiHuman-Testbench (MHTB) [4] with 1,800 samples and 5550 faces, uniformly spanning 1-5 people per sample; and Multi-ID-Test [38] with 1-4 faces biased toward 1-2 person scenes.
Evaluation Metrics. Following the setup in MHTB [4], we employ four complementary metrics: Count Accuracy for correct number of generated people; Hungarian ID Similarity evaluates identity preservation using facial similarity with Hungarian matching to establish and quantify optimal face correspondences; HPSv2 measures overall image quality and prompt alignment; and Action-S/C probes action correctness separately for simple and complex actions, using Gemini Flash as the MLLM. We also report the Unified Metric [4] which balances identity preservation and prompt alignment.